OpenStudy (anonymous):

a right triangle has a hypotenuse of 9 inches and a leg of 4 inches. find the length of the other leg. round to nearest thousandth

ganeshie8 (ganeshie8):

oh u figured out the answer already !

OpenStudy (anonymous):

no

ganeshie8 (ganeshie8):

okay.. i thought u figured out the solution as u closed the q

ganeshie8 (ganeshie8):

so its a right triangle

OpenStudy (anonymous):

i was trying to figure it out on my own and it didnt work.

OpenStudy (anonymous):

a right triangle has a hypotenuse of 9 inches and a leg of 4 inches. find the length of the other leg. round to nearest thousandth

ganeshie8 (ganeshie8):

hypotenuse and one leg are given

ganeshie8 (ganeshie8):

you need to find the other leg, right ?

OpenStudy (anonymous):

right

ganeshie8 (ganeshie8):

its easy, we have a formula for right triangle. hypotenuse^2 = leg1^2 + leg2^2

ganeshie8 (ganeshie8):

9^2 = 4^2 + leg2^2 => 81 = 16 + leg2^2

ganeshie8 (ganeshie8):

you can solve leg2 now ?

OpenStudy (anonymous):

64???

ganeshie8 (ganeshie8):

wrong... try step by step

ganeshie8 (ganeshie8):

81 = 16 + leg2^2 subtract 16 both sides => 81-16 = leg2^2 => 65 = leg2^2

OpenStudy (anonymous):

im sooo burnt out on numbers

ganeshie8 (ganeshie8):

its okay... more u practice... the more u feel confy :D

ganeshie8 (ganeshie8):

leg2^2 = 65 => leg2 = \(\sqrt{65}\)

ganeshie8 (ganeshie8):

put it in calculator and see what u get

OpenStudy (anonymous):

8.062257748298

ganeshie8 (ganeshie8):

now, question is asking to round this to nearest thousandth

ganeshie8 (ganeshie8):

u familiar with rounding thingy

ganeshie8 (ganeshie8):

?

OpenStudy (anonymous):

not really but ill try.

OpenStudy (anonymous):

8.1?

ganeshie8 (ganeshie8):

thats correct to tenth

ganeshie8 (ganeshie8):

go two more digits to left, and then round !

ganeshie8 (ganeshie8):